OpenStudy (anonymous):

\[F=(ye^{xy}\ln(1+z^2))i + (xe^{xy}\ln(1+z^2)+1)j + (\frac{2ze^{xy}}{1+z^2})k\] The potential function of F, I worked out to be: \[f(x,y,z)=e^{xy}\ln(z^2+1)+y+c\] \[G = (ye^{xy}\ln(1+z^2)+x)i + (xe^{xy}\ln(1+z^2)+1)j + (\frac{2ze^{xy}}{1+z^2})k\] Let C be the curve x=t, y=0, z=t(1-t) with \[0\le t \le 1\] which connects \[x_{0}=(0,0,0)\] with \[x_{1}=(1,0,0)\] Find \[\int\limits_C G.dr\]

OpenStudy (anonymous):

@sazap10 , Would you mind typing only the question separately because I can't distinguish between what you have done and what is given in the question

OpenStudy (anonymous):

Let \[G = (ye^{xy}\ln(1+z^2)+x)i + (xe^{xy}\ln(1+z^2)+1)j + (\frac{2ze^{xy}}{1+z^2})k\] Let C be a curve of the form x=t, y=0, z=t(1-t) with \[0\le t \le 1\] which connects \[x_{0}=(0,0,0)\] with \[x_{1}=(1,0,0)\] Find \[\int\limits_C G.dr\]

OpenStudy (anonymous):

@experimentX will help you in this :)

OpenStudy (experimentx):

what is G??

OpenStudy (anonymous):

If I am not wrong , G is flux ?? Am I right??

OpenStudy (anonymous):

a vector field i think

OpenStudy (experimentx):

Change all x, y, and z to t's take a dot product .. and integrate.

OpenStudy (experimentx):

dr = idx + jdy + kdz

OpenStudy (anonymous):

i think u are supposed to use the potential function f

OpenStudy (anonymous):

but i idk how since G is different from F

OpenStudy (experimentx):

F and G are same ... expect for the fact that there is extra x on i component of G

OpenStudy (anonymous):

then could i use the potential function f to find the circulation? i did that and got the answer as 0. Which doesnt seem right to me

OpenStudy (experimentx):

it's only changing around x axis ... i don't think these two points are equipotential surface.

OpenStudy (experimentx):

Change all x,y,z to t's and integrate \[ \int_{0}^{1}(ye^{xy}\ln(1+z^2)+x)dx\]

OpenStudy (anonymous):

G is a conservative field, then the work done from point A to point B is fB) - f(A) where f is the potential of the conservative Field G

OpenStudy (anonymous):

The answer is f(1,0,0)-f(0,0,0) =ln(1)